OpenStudy (anonymous):

How does a irrational power multiply your function? For example x^(3/2) and x^(6/7) How can I write it down without square roots? Because I need to write down (1+sqrt(x))^(3/2)

OpenStudy (anonymous):

\[\left( 1+\sqrt{x} \right)^{\frac{ 3 }{ 2 }}=\sqrt{\left( 1+\sqrt{x} \right)^3}\]